OpenStudy (anonymous):

Determine the domain, range, and vertical asymptote f(x)=log(5)(x+4) CAN SOME ONE PLEASE HELP????????

OpenStudy (anonymous):

Is that \[f(x)=log_5{(x+4)}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay we know that we can't take the log(0) because it is undefined. So the value for x where this occurs is our vertical asymptote. x+4=0 x=-4 ->vertical asymptote occurs at -4 Now since we also can't have a log(-x) . The x values have to be greater that -4. \[Domain: x \in \mathbb{R}:x > -4\] The range can be all real numbers (R). \[Range: \mathbb{R}\]